# Adaptive Group Robust Ensemble Knowledge Distillation

### v1

@@ 4 Adaptive Group Robust Ensemble Knowledge Distillation (AGRE-KD)
Figure 2: Overview of AGRE-KD. Detailed algorithm is provided in Supplementary A.
In this work, we considered each teacher in the ensemble to have the same architecture and trained using different random initializations. Prior work has shown ensemble models with different random initializations are diverse enough to improve the performances (Ganaie et al.,, 2022). This work shows that while deep ensemble models can improve the worst-case group (Ko et al.,, 2023), it is not necessarily the case in ensemble knowledge distillation. To address this problem, we propose AGRE-KD, an adaptive ensembling knowledge distillation strategy that ensures the student model captures robust knowledge from the teachers. Figure 2 provides an overview of our proposed method. AGRE-KD relies on a model pretrained with ERM that captured the dataset’s spurious correlation (biased). Intuitively, suppose a student model takes gradient steps toward the direction that minimizes its KD loss with the biased model. In that case, the resulting student model will likely capture and even amplify the reliance on the spurious correlation, i.e., the local/global minimum in that direction likely provides the worst performance for the underrepresented subgroups. Note that an additional biased model does not add extra complexity compared to related work using a reference classifier to identify the worst-performing subgroups (Nam et al.,, 2020; Liu et al.,, 2021; Kenfack et al., 2024b, ). Furthermore, as we will see in the experiments, our proposed method works with any biased model trained using ERM without further modifications.
@@
[…]
@@
And we train the student model with the following final loss $\mathcal{L}=\alpha\mathcal{L}_{\text{wKD}}+(1-\alpha)\mathcal{L}_{\text{cls}}$
In this work, we focused on unsupervised knowledge distillation, i.e., we set $\alpha=1$. This means we train the student model only using the output of the teacher models that are eventually debiased. Supplementary A provides a detailed algorithm of the training process of AGRE-KD.
@@ 5 Experimental Results
In this section, we present the experimental setup and the empirical results. We compare our adaptive ensemble knowledge distillation to other methods and demonstrate our method’s effectiveness in improving the student model’s worst-case performance.
@@ 5.1 Setup
We evaluate the worst-case performance of the proposed method on three classification tasks: two from the vision domain (Waterbird and CelebA dataset) and one from the language domain (CivilComments)
  • Waterbird (Sagawa et al.,, 2019; Liu et al.,, 2021) is a dataset of birds derived from Caltech-UCSD Birds (CUB) (Wah et al.,, 2011) by synthetically creating a spurious correlation between bird species and the background. In particular, the class label is the type of bird appearing in the image (waterbirds, landbirds), and the background landscape (water, land) spuriously correlates with the bird type. Here, the minority subgroups represent images with the background landscape not aligned with the bird type, i.e., {waterbird, land background} and {landbird, water background}.
  • CelebA (Liu et al.,, 2015) dataset contains images of celebrities with 40 facial attributes. In this dataset, the attribute hair color is spuriously correlated gender. We consider hair color {blond, non-blond} as the class label and gender {male, female} as group information.
  • CivilComments (Koh et al.,, 2021) is a textual dataset collected from online comments. The task is to predict whether a comment is toxic or non-toxic. The label is spuriously correlated with comments related to some demographic subgroups such as gender (male, female), race (white, black), and sexual orientation (LGBT). We consider a binary indicator of comments related to these demographic subgroups as spurious group information.
@@
Following prior work (Tiwari et al.,, 2024; Lee and Lee,, 2023), we use the Resnet-18 (He et al.,, 2016) architecture for the student model and the Resnet-50 (He et al.,, 2016) architecture for the teacher models for the forvision tasks. Both networks are pretrained on the ImageNet-1K (Russakovsky et al.,, 2015) dataset. For the language task, we used the BERT (Devlin et al.,, 2019) model for teachers and the DistilBERT (Sanh et al.,, 2019) for the student model; and language models are pretrained on Book Corpus and English Wikipedia. Following related works on KD (Du et al.,, 2020; Fukuda et al.,, 2017; Chen et al.,, 2022), we set the temperature hyperparameter $\tau=4$ (Eq. 1) and show in an ablation study in Supplementary C (Figure 5) that increasing $\tau$ can exert positive effect on WGA up to certain values. We provide further details about hyperparameters in the Supplementary B.
Teacher training.
We train each teacher model in the ensemble independently using standard ERM and cross-entropy loss with the ground truth class labels. Teacher models have the same architecture and hyperparameters and only differ in random seeds used for weights initialization; prior studies have shown that independent training with different random weight initializations can provide models sufficient diversity to improve the performance of the model ensemble (Allen-Zhu and Li,, 2020; Ganaie et al.,, 2022). As aforementioned, we obtain debiased teachers using the approach introduced by Kirichenko et al., (2022) to retrain the last layer of the ERM model on the small proportion of the held-out group-balanced dataset. We perform the retraining step of the last layer using group-balanced batch sampling instead of the averaging models trained over group-balanced subsets of the data (LaBonte et al.,, 2024). This debiasing process is very simple and computationally inexpensive since it involves training a logistic regression model on a smaller dataset.
For the biased model used by our method to compute teachers’s weights, we randomly select one teacher model trained with ERM and without DFR. We provide in Supplementary C (Table 3) experiment with different choices of biased model showing the robustness of the proposed gradient-based weighting to the biased model choice. Following Kirichenko et al., (2022); LaBonte et al., (2024), we use half of the validation set of each benchmark to perform last-layer retraining with DFR; as we do not use the group and class labels during the KD training, we do not perform further hyperparameter tuning or model selection.
Baselines
@@
  • Deep Ensemble: This baseline corresponds to deep ensembling using a majority voting scheme of models with the same capacity as the student model. In particular, given a set of models, the predicted class label represents the class that received the most votes for models in the ensemble.
  • KD with averaged teachers outputs (AVER) (You et al.,, 2017): Here, we perform standard knowledge distillation following equation 3 by minimizing the KL loss between the student’s softmax output and the averaged softened outputs (dark knowledge) from teachers.
  • Random (Fukuda et al.,, 2017): During each mini-batch training, this method randomly selects a teacher model from the ensemble to train the student model.  Fukuda et al., (2017) referred to this technique as switched-training as the weights of the student model are updated by switching across teacher labels at the minibatch level.
  • AE-KD (Du et al.,, 2020): This method is an adaptive ensembling distillation technique closest to ours. However, the method postulates that when teachers have conflicting gradient directions, a multi-objective optimization problem is solved to select the gradient direction (teachers) that satisfies most of the teachers in the ensemble.

### v2

@@ 4 Adaptive Group Robust Ensemble Knowledge Distillation (AGRE-KD)
Figure 2: Overview of AGRE-KD.
In this work, we considered each teacher in the ensemble to have the same architecture and trained using different random initializations. Prior work has shown ensemble models with different random initializations are diverse enough to improve the performance (Ganaie et al., 2022). The first contribution of this work is showing that, unlike deep ensemble models that can improve the WGA (Ko et al., 2023), ensemble knowledge distillation tends to amplify bias in the distilled model.
To address this problem, we propose AGRE-KD, an adaptive ensembling knowledge distillation strategy that ensures the student model captures robust knowledge from the teachers. AGRE-KD relies on a model pretrained with ERM that captured the dataset’s spurious correlation (biased model). Intuitively, suppose a student model takes gradient steps toward the direction that minimizes its KD loss with the biased model. In that case, the resulting student model will likely capture and even amplify the reliance on the spurious correlation, i.e., the local/global minimum in that direction likely provides the worst performance for the underrepresented subgroups. Following this intuition, we derive a weighting scheme that prioritizes teacher models whose gradient direction disagrees with the biased ERM model. Figure 2 provides an overview of our proposed method.
Teacher training.
We train each teacher model in the ensemble independently using standard ERM and cross-entropy loss with the ground truth class labels. Teacher models have the same architecture and hyperparameters and only differ in random seeds used for weights initialization; prior studies have shown that independent training with different random weight initializations can provide sufficient ensemble diversity to improve the performance of the model ensemble (Allen-Zhu & Li, 2020; Ganaie et al., 2022). As aforementioned, we obtain debiased teachers using the approach introduced by Kirichenko et al. (2022) to retrain the last layer of the ERM model on a small proportion of the held-out group-balanced dataset. We perform the retraining step of the last layer using group-balanced batch sampling instead of the averaging models trained over group-balanced subsets of the data (LaBonte et al., 2024). This debiasing process is very simple and computationally inexpensive since it involves training a logistic regression model on a smaller dataset. Following Kirichenko et al. (2022); LaBonte et al. (2024), we use half of the validation set of each benchmark to perform last-layer retraining with DFR; as we do not use the group and class labels during the KD training, we do not perform further hyperparameter tuning or model selection. After training, teacher models are frozen, and their output is aggregated to train student models as illustrated in figure 2.
Biased model.
For the biased model used by our method to compute teachers’ weights, we randomly select one teacher model trained with ERM. ERM-trained models are known to be sensitive to spurious correlation (Kirichenko et al., 2022; Ye et al., 2024). This provides us with intuition that the gradient direction updates of a model during ERM training are biased towards the majority, and leverage this intuition to upweight teacher models that deviate from the direction provided by an ERM model. Moreover, as we show via an ablation study in Supplementary D (Table 5), the performance of our gradient-based weighting scheme remains robust to different choices of ERM model architectures.
@@
For a given minibatch, we compute the student KD loss regarding the biased model $b$ and each teacher $t$. Before aggregating the teachers’ outputs, we compute sample weights based on the similarity between the gradient direction of each teacher and the biased model. Consider $\ell^{t}_{i}(\theta)$ the KD loss on sample $i$ regarding the $t$-teacher, and $\ell^{b}_{i}(\theta)$ KD loss regarding the biased model, with $\theta$ the parameters of the student model. The dot product ($\langle\cdot,\cdot\rangle$) between the normalized [^2] Table 6 in the Supplementary shows the importance of ignoring the magnitude of the gradients. gradients of the student KD loss with the teacher and the biased model indicates which teachers align with the biased model in gradient space. In particular, when $\langle\nabla\ell^{t}_{i}(\theta),\nabla\ell^{b}_{i}(\theta)\rangle>0$ both models have the same gradient directions and their gradients are in the opposite direction when $\langle\nabla\ell^{t}_{i}(\theta),\nabla\ell^{b}_{i}(\theta)\rangle<0$. In extreme cases, when the dot product of the gradients is closer to $1$, both models have precisely the same directions; if the dot product is close to 1-1, the models are in opposite directions. Thus, to penalize teachers with a gradient direction closer to the biased model, we downweight teachers’ output for samples having dot products get closer to $1$ and upweight samples as their dot products get closer to 1-1. More formally, we compute the sample-wise teacher’s weight as follows:
@@
And we train the student model with the following final loss
$\mathcal{L}=\alpha\mathcal{L}_{\text{wKD}}+(1-\alpha)\mathcal{L}_{\text{cls}}$ (7)
As aforementioned, we focus on unsupervised knowledge distillation, i.e., we set $\alpha=1$ to investigate the robustness of the "dark knowledge" provided by an ensemble of eventually debiased teachers. Supplementary A provides a detailed algorithm of the training process of AGRE-KD. Note that the use of an additional biased model to orchestrate the distillation process does not add extra complexity compared to related work using a reference classifier to identify the worst-performing subgroups (Nam et al., 2020; Liu et al., 2021; Kenfack et al., 2024a). Furthermore, as we will see in the experiments, our proposed method is robust to the choice of biased model architecture trained using ERM without.
@@ 5 Experimental Results
In this section, we present the experimental setup and the empirical results. We experiment with several ensemble KD techniques and analyze their impact on the worst-group accuracy of the student model. We compare our adaptive ensemble knowledge distillation to these methods and demonstrate its effectiveness in improving the student model’s worst-group performance.
@@ 5.1 Setup
We evaluate the worst-case performance of the proposed method on four classification tasks: one synthetic dataset (Colored MNIST), two real-world datasets from the vision domain (Waterbirds and CelebA datasets), and one real-world dataset from the language domain (CivilComments).
  • Colored MNIST following Nam et al. (2020); Li & Vasconcelos (2019), we alter the MNIST dataset (LeCun et al., 2002), designed for digit classification, by artificially adding a Color attribute. To define the Color, we randomly select ten distinct RGB values and fix them across all experiments. Each of these values serves as the mean of a 3-dimensional Gaussian distribution, forming ten distinct Color distributions. To create correlations between digits and colors, we pair each digit with one of these distributions. A bias-aligned sample is generated by coloring a digit with an RGB value drawn from its paired distribution, while a bias-conflicting sample is produced by sampling from any of the remaining nine color distributions (Nam et al., 2020). Under this setup, a model that captures the spurious correlation will predict color instead of the digits.
  • Waterbirds (Sagawa et al., 2019; Liu et al., 2021) is a dataset of birds derived from Caltech-UCSD Birds (CUB) (Wah et al., 2011) by synthetically creating a spurious correlation between bird species and the background. In particular, the class label is the type of bird appearing in the image (waterbirds, landbirds), and the background landscape (water, land) spuriously correlates with the bird type. Here, the minority subgroups represent images with the background landscape not aligned with the bird type, i.e., {waterbird, land background} and {landbird, water background}.
  • CelebA (Liu et al., 2015) dataset contains images of celebrities with 40 facial attributes. In this dataset, the attribute hair color is spuriously correlated with gender. We consider hair color {blond, non-blond} as the class label and gender {male, female} as group information.
  • CivilComments (Koh et al., 2021) is a textual dataset collected from online comments. The task is to predict whether a comment is toxic or non-toxic. The label is spuriously correlated with comments related to some demographic subgroups such as gender (male, female), race (white, black), and sexual orientation (LGBT). We consider a binary indicator of comments related to these demographic subgroups as spurious group information.
@@
For the Colored MNIST dataset, we use MobileNetV2 (Sandler et al., 2018) for the teacher models and a simple CNN with two convolutional layers as for the student model. For the real-world vision tasks (CelebA and WaterBirds), following prior work (Tiwari et al., 2024; Lee & Lee, 2023), we use the Resnet-18 (He et al., 2016) architecture for the student model and the Resnet-50 (He et al., 2016) architecture for the teacher models. Both networks are pretrained on the ImageNet-1K (Russakovsky et al., 2015) dataset. For the language task, we use the BERT (Devlin et al., 2019) model for teachers and the DistilBERT (Sanh et al., 2019) for the student model; and language models are pretrained on Book Corpus and English Wikipedia. Following related works on KD (Du et al., 2020; Fukuda et al., 2017; Chen et al., 2022), we set the temperature hyperparameter $\tau=4$ (Eq. 1) and show in an ablation study in Supplementary D (Figure 5) that increasing $\tau$ can exert positive effect on WGA up to certain values. We provide further details about hyperparameters in the Supplementary B.
Baselines.
@@
  • Deep Ensemble (Deep Ens.): This baseline corresponds to deep ensembling using a majority voting scheme of models with the same capacity as the student model. In particular, given a set of models, the predicted class label represents the class that received the most votes for models in the ensemble.
  • One-hot: A student model trained with ground truth labels without knowledge distillation.
  • Random (Fukuda et al., 2017): During each mini-batch training, this method randomly selects a teacher model from the ensemble to train the student model.  Fukuda et al. (2017) referred to this technique as switched-training as the weights of the student model are updated by switching across teacher labels at the minibatch level.
  • AVER (KD with averaged teachers’ outputs (You et al., 2017)): Here, we perform standard knowledge distillation following equation 3 by minimizing the KL loss between the student’s softmax output and the averaged softened outputs (dark knowledge) from teachers.
  • AE-KD (Du et al., 2020): This method is an adaptive ensembling distillation technique closest to ours. However, the method postulates that when teachers have conflicting gradient directions, a multi-objective optimization problem is solved to select the gradient direction (teachers) that satisfies most of the teachers in the ensemble.
@@ 5.2 Results
We train each model using three random seeds and report the means and standard deviations. We consider ensemble distillation with ten teacher models randomly sampled from a poll of pretrained (biased) teachers across independent random seeds. In addition to the teacher’s performance and deep ensemble approach, we report the performance of the "student" trained only using the ground truth class labels (One-hot), i.e., without knowledge distillation. We report the average accuracy and WGA. We consider models trained using ERM or with last-layer retraining for debiasing. Specifically, for ensemble knowledge distillation methods, we train the student model with biased teachers (ERM model) and debiased teachers (last layer retrained with DFR (Kirichenko et al., 2022)).
5.2.1 Results on the synthetic dataset
We considered different settings of the Colored MNIST (CMNIST) dataset by varying the ratios of bias-aligned samples in the training dataset, i.e., the proportion of samples where the color and digit correspond $\{99.5\%,99\%,98\%,\text{and}\;95\%\}$. This means for configuration with a ratio of 99.5%, only 0.005% of samples will have a digit-color mismatch, and decreasing the ratio reduces the strength of the spurious correlation in the training dataset. We train every baseline under each bias ratio configuration and report the average and worst-group test accuracy.
Table 1 presents the performance of different approaches on Colored MNIST under varying levels of spurious correlation. Several important observations can be drawn.
The teacher models illustrate the critical role of debiasing. Without debiasing, teachers perform poorly on the minority group (WGA as low as 30.5% under 99.5% spurious correlation), even average accuracy appears lower (37.4%) since the test set mainly contains samples where spurious correlation does not apply. Once debiased, however, teacher models show clear improvements in WGA across all regimes, confirming that debiased supervision is key to counteracting extreme spurious correlations.
Deep Ensembles models already provide a substantial performance boost: simply averaging biased models significantly boosts robustness compared to a single teacher, leading to WGA improvements of  20–40 points in extreme regimes (e.g., 29.2% to 69.8% at 99.5%). Interestingly, when ensemble predictions are used for knowledge distillation (AVER, AEKD, or AGRE-KD with biased teachers), performance patterns diverge. In highly biased settings (CMNIST-99.5%, CMNIST-99%), ensemble KD tends to underperform relative to the raw Deep Ensemble, often worsening WGA (e.g., AVER and AEKD around 49–54% WGA vs. 69.8% for Deep Ens. in CMNIST-99.5%). However, in less extreme regimes (CMNIST-98% and CMNIST-95%), ensemble KD begins to catch up and even surpass the teacher ensemble, suggesting that distillation stabilizes and refines knowledge transfer when correlations are milder.
Both AVER and AEKD, which attempt to leverage ensemble outputs in different ways, offer moderate improvements over naïve or random strategies. For example, AEKD achieves 66.3% WGA at 99% with a debiased teacher, substantially higher than One-hot or Random. However, their effectiveness remains inconsistent: in extreme regimes, they still fail to match the robustness of the Deep Ensemble, and in milder regimes, they are often outperformed by stronger approaches.
Across all bias regimes, AGRE-KD consistently yields the best performance when combined with debiased teachers. It achieves the highest WGA at every correlation level, from 59.9% under the most extreme bias (CMNIST-99.5%) to 88.7% in CMNIST-95%. Notably, AGRE-KD is the only KD method that not only closes the gap with the raw Deep Ensemble in highly biased regimes but surpasses it in less extreme regimes. Even with biased teachers, AGRE-KD matches or exceeds other KD baselines, demonstrating its robustness to spurious correlation during the distillation process.
Models Debiased KD CMNIST-99.5% CMNIST-99% CMNIST-98% CMNIST-95%
Average WGA Average WGA Average WGA Average WGA
Teacher _ $37.4_{\pm 0.69}$ $30.5_{\pm 0.61}$ $55.3_{\pm 0.92}$ $50.4_{\pm 1.00}$ $71.7_{\pm 0.45}$ $68.7_{\pm 0.56}$ $87.7_{\pm 0.10}$ $86.3_{\pm 0.14}$
_ $58.4_{\pm 1.84}$ $54.7_{\pm 2.31}$ $66.0_{\pm 0.97}$ $63.1_{\pm 1.00}$ $80.3_{\pm 0.60}$ $78.1_{\pm 0.70}$ $83.4_{\pm 0.53}$ $81.7_{\pm 0.58}$
Deep Ens. _ $36.2_{\pm 0.38}$ $29.2_{\pm 0.44}$ $56.2_{\pm 0.25}$ $51.4_{\pm 0.22}$ $74.4_{\pm 0.09}$ $71.5_{\pm 0.15}$ $9.6_{\pm 0.09} 88$$8.4_{\pm 0.12}$
_ $72.8_{\pm 0.27}$ $69.8_{\pm 0.23}$ $77.7_{\pm 0.11}$ $75.3_{\pm 0.09}$ $84.2_{\pm 0.16}$ $82.5_{\pm 0.19}$ $90.0_{\pm 0.09}$ $88.9_{\pm 0.09}$
One-hot $26.9_{\pm 6.52}$ $19.0_{\pm 7.44}$ $29.3_{\pm 6.94}$ $21.6_{\pm 7.93}$ $34.5_{\pm 7.93}$ $27.4_{\pm 9.01}$ $47.0_{\pm 7.91}$ $41.4_{\pm 8.97}$
$53.6_{\pm 6.29}$ $52.6_{\pm 5.91}$ $51.9_{\pm 5.68}$ $51.7_{\pm 5.63}$ $81.5_{\pm 1.58}$ $80.0_{\pm 1.82}$ $61.8_{\pm 4.83}$ $60.9_{\pm 4.37}$
Random $25.0_{\pm 2.41}$ $16.7_{\pm 2.58}$ $26.5_{\pm 3.31}$ $18.5_{\pm 3.58}$ $29.4_{\pm 4.24}$ $21.7_{\pm 4.61}$ $41.1_{\pm 3.92}$ $34.6_{\pm 4.29}$
$53.2_{\pm 5.35}$ $48.5_{\pm 6.02}$ $9.7_{\pm 5.90} 44$$4.5_{\pm 6.49}$ $39.3_{\pm 4.07}$ $32.6_{\pm 4.45}$ $41.8_{\pm 4.67}$ $35.5_{\pm 5.12}$
AVER $33.9_{\pm 0.48}$ $26.5_{\pm 0.51}$ $51.6_{\pm 0.14}$ $46.2_{\pm 0.21}$ $72.2_{\pm 0.59}$ $69.1_{\pm 0.66}$ $9.3_{\pm 0.47} 88$$8.1_{\pm 0.54}$
$54.1_{\pm 0.23}$ $49.0_{\pm 0.28}$ $62.8_{\pm 0.42}$ $58.7_{\pm 0.48}$ $86.5_{\pm 0.55}$ $5.0_{\pm 0.63} 88$$8.5_{\pm 0.19}$ $87.3_{\pm 0.22}$
AEKD $36.2_{\pm 0.76}$ $\underline{29.2}_{\pm 0.81}$ $55.0_{\pm 0.83}$ $50.0_{\pm 0.84}$ $73.4_{\pm 0.78}$ $70.5_{\pm 0.81}$ $9.7_{\pm 0.09} 88$$8.6_{\pm 0.07}$
$58.8_{\pm 0.80}$ $54.5_{\pm 1.29}$ $9.4_{\pm 1.56} 66$$6.3_{\pm 1.65}$ $86.6_{\pm 0.65}$ $5.1_{\pm 0.68} 88$$8.1_{\pm 0.45}$ $86.9_{\pm 0.53}$
AGRE-KD $\underline{36.3}_{\pm 0.65}$ $\underline{29.2}_{\pm 0.65}$ $\underline{55.5}_{\pm 0.56}$ $\underline{51.5}_{\pm 0.59}$ $\underline{74.2}_{\pm 0.21}$ $\underline{71.4}_{\pm 0.26}$ $\underline{90.2}_{\pm 0.31}$ $\underline{89.1}_{\pm 0.34}$
$\textbf{63.9}_{\pm 0.33}$ $\textbf{59.9}_{\pm 0.37}$ $\textbf{70.7}_{\pm 0.43}$ $\textbf{67.5}_{\pm 0.53}$ $\textbf{87.5}_{\pm 0.51}$ $\textbf{86.2}_{\pm 0.57}$ $\textbf{89.0}_{\pm 0.48}$ $\textbf{88.7}_{\pm 0.49}$
Table 1: Results on the Colored MNIST dataset. We report the average and worst-group test accuracy (WGA) for different ratios of bias-aligned samples ($\{99.5\%,99\%,98\%,\text{and}\;95\%\}$). The Debiased column indicates whether the model involves debiasing with DFR or whether teacher models are debiased when the KD column is checked (✓). Bolded represents the best-performing student with the debiased teacher ensemble and underlined represents the best-performing with the biased teacher ensemble.
5.2.2 Results on real-world datasets
Table 2 summarizes the main results of the paper on real-world datasets, from which we draw the following observations:
  • Ensemble knowledge distillation can amplify bias. From Table 2, we first observe that deep ensembles improve worst-group accuracy (WGA) relative to single models (one-hot). For example, on Waterbirds, WGA increases from $54.1\%$ (biased model) to $59.3\%$ (biased deep ensemble) and from $86.7\%$ (debiased model) to $90.0\%$ (debiased deep ensemble). On CelebA, the jump is even larger: WGA improves from $32.3\%$ (biased model) to $37.2\%$ (biased deep ensemble) and from $88.9\%$ (debiased teacher) to $90.3\%$ (debiased deep ensemble). This confirms that ensembling stabilizes predictions and mitigates reliance on spurious features.
    However, when ensemble predictions are used for knowledge distillation, the picture changes. In highly biased regimes, ensemble KD often reduces WGA. For instance, on Waterbirds, Random KD drops WGA to $39.6\%$, nearly $20$ points below the biased deep ensemble ($59.3\%$). Similarly, AVER KD achieves only $46.4\%$ WGA on Waterbirds and $28.8\%$ on CelebA, representing losses of more than $10$ points compared to their ensemble teachers. These results suggest that switching between teachers or naively averaging teachers’ output during distillation amplifies bias, particularly harming minority groups, even if average accuracy remains high (e.g., CelebA average accuracy stays around $95.5\%$). When comparing to the one-hot baseline, we observe mixed results, but generally, ensemble KD can increase average accuracy compared to the "One-hot" baseline, but they are also more prone to bias amplification. This reinforces our finding that while ensemble KD offers advantages, debiasing mechanisms are essential to make it competitive with simpler approaches like direct training or DFR.
  • DFR teachers lead to robust student models. When we train the teacher models using ERM and then retrain the classification layer with the group-balanced set, the resulting student models achieve significantly better worst-case group accuracy compared to those trained only with ERM teacher models. We only train the student model using the aggregated teachers’ outputs without any feature distillation. For example, on CelebA, Random KD training with debiased teachers improves WGA from $27.2\%$ (biased) to $85.1\%$ (debiased), a remarkable +57.9 point gain. Similar patterns hold for the other ensemble KD methods with AVER KD ($28.8\%\to 83.4\%$), AE-KD ($46.9\%\to 85.0\%$), and AGRE-KD ($55.0\%\to 87.9\%$).
    These results demonstrate that the retrained classification layer can provide pseudo-label distribution (dark knowledge) that reduces students’ reliance on spurious features. This shows that by mimicking the teacher’s outputs, the classifier layer of the student model also downweights the spurious features in its last layers (Izmailov et al., 2022). As the temperature parameter ($\tau$) also alters the teacher’s output by softening its predicted probability distribution, we provide an ablation study (Appendix C.2) showing that increasing $\tau$ positively impacts the worst-group accuracy. We also show in Appendix LABEL:app:ablation_alpha experiments with different values of $\alpha$ (Eq. 37), controlling the balance between updating the student model with ground truth label and ensemble knowledge distillation loss. The results show that the student model gets less biased as more weight is given to the knowledge distillation loss. For example, on the Waterbirds dataset, the WGA increases linearly by up to $25\%$ with increased values of $\alpha$.
    On the other hand, the improved students’ WGA across ensemble distillation methods still does not match the WGA of the debiased teachers or debiased deep ensemble. For example, on Waterbirds, AVER KD with debiased teachers achieves $82.9\%$ WGA, $7\%$ below the $90.9\%$ WGA of the debiased teacher. We attribute this to the smaller capacity of the student model, which we discuss in the next experiment.
  • Our weighting scheme consistently improves WGA. Finally, our proposed AGRE-KD achieves the most robust results across datasets. With debiased teachers, AGRE-KD attains the highest WGA in all three benchmarks: $87.9\%$ on Waterbirds, $91.9\%$ on CelebA, and $75.9\%$ on CivilComments. This represents improvements of $2.9$, $3.0$, and $0.9\%$, respectively, over the best competing KD methods (AEKD or AVER). Importantly, on CelebA, AGRE-KD even outperforms the deep ensemble itself ($91.9\%$ vs. $90.3\%$ WGA), showing that our weighting scheme extracts more bias-free knowledge than naive ensembling. When all teachers are biased, AGRE-KD still identifies and upweights the least biased members. For example, on Waterbirds, it achieves $55.0\%$ WGA, outperforming AVER KD($46.4\%$) and Random KD ($39.6\%$). On CelebA, it reaches $37.6\%$, slightly higher than the biased teacher baseline ($37.5\%$). While the improvements in this setting are modest—since all teachers share similar spurious gradients—the method still avoids the large collapses observed with other KD strategies. In the Supplementary D (Table 7), we provide the group-wise accuracy of each baseline method on the Waterbirds and CelebA datasets, showing how each subgroup is impacted by each method.